OpenStudy (anonymous):

The distance between the points S(x,5) and U(-4, 13) is 10. What is the value of X?

OpenStudy (anonymous):

\[\sqrt{(x+4)^2+(5-13)^2}=10 \implies x^2+8x+16+64=100\] \[\implies x^2+8x-20=0 \implies (x+10)(x-2)=0 \implies x=2, x=-10\]